SwaggyMark:

At LaGuardia Airport for a certain nightly flight, the probability that it will rain is 0.09 and the probability that the flight will be delayed is 0.06. The probability that it will rain and the flight will be delayed is 0.05. What is the probability that the flight would be delayed when it is raining? Round your answer to the nearest thousandth.

WickedHeart:

Since at LaGuardia Airport, for a certain nightly flight, the probability that it will rain is 0.19 and the probability that the flight will be delayed is 0.15, while the probability that it will not rain and the flight will leave on time is 0.74 , to determine what is the probability that the flight would be delayed when it is not raining, the following calculation must be performed: Probability that it will not rain = 1 - probability that it will rain X = 1 - 0.19 X = 0.81 Probability that the flight would be delayed when it is not raining = probability that it is not raining x probability that the flight will be delayed \[X = 0.81 \times 0.15\] \[X = 0.1215\] \[0.1215 \times 100 = ?\]

WickedHeart:

Now what would the Answer be?

ExclusiveKylaa:

A, The chances of it raining are 0.09, while the chances of the flight being delayed are 0.18.

WickedHeart:

@exclusivekylaa wrote:
A, The chances of it raining are 0.09, while the chances of the flight being delayed are 0.18.
Direct ANswers are not allowed thats why i had the person work it out